OpenStudy (anonymous):

Solve for x. Show the work 81^(4x-2)=3

OpenStudy (anonymous):

\[81^{4x-2}=3\] \[3^{4(4x-2)}=3\] \[4(4x-2)=3\] you can do now yourself